[Hallc_running] [New Logentry] Hall C update (Dec. 9)

gaskelld at jlab.org gaskelld at jlab.org
Sat Dec 9 13:15:02 EST 2017

Logentry Text:
Last night, beam was successfully delivered to Hall C. After the ion chamber calibration procedure, Jay Benesch's SHMS steering test plan was performed. While the large downstream BPMs were not functioning perfectly, we were able to observe some change in the beam trajectory when ramping up the SHMS to 2.2 GeV at 7.5 degrees, with the steering effects largely gone by the time the SHMS was rotated to 8.5 degrees.

We ran into some trouble with the harps at 3H07A and 3H07B this morning, but hopefully this has been fixed with a reboot of the ioc.

Right now, the machine is down - initially to do some PSS work to bring up Hall A, but along the way ran into another problem in the South Linac.

After beam restoration, we will check the carbon hole target and attempt the BPM/harp calibration again.  Then we will move on to the SHMS/HMS detector and optics commissioning program.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3501627
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20171209/ed52865f/attachment.html>

More information about the Hallc_running mailing list